Estimate a = 0 the area to = 6 under the wire (x) = 2xt rom. using n = 3 right Vendpoint sudargles.
Answer #1


fed -2X + 13 = +16) - Area using right fler and point is, A = $x[419) + f(9) + HC)] 97409 Your bone = 900 712=27 5+® -A=275+2